OpenStudy (rissa):

Suppose you had d dollars in your bank account. You spent $22 but have at least $28 left. How much money did you have initially? Write and solve an inequality that represents this situation.

OpenStudy (anonymous):

50

OpenStudy (anonymous):

oohh wait..

OpenStudy (lilg132):

x = 22 + 28 22 + 28 = 50

OpenStudy (anonymous):

28+22=50 d=50

OpenStudy (anonymous):

yea there u go :)

OpenStudy (lilg132):

sorry instead of x use d

OpenStudy (lilg132):

cough